OpenStudy (anonymous):

Jake is now one-quarter of his father’s age and in 7 years time his age will be one - third of his father. How old is Jake now?

OpenStudy (anonymous):

J=F/4 J+7=F/3 Sun J into equation 2 and solve. Here J and F are the ages of Jake and his father

OpenStudy (anonymous):

its confusing...

OpenStudy (anonymous):

J = 7 / (1/3) / (1/4) = ?